Suppose you could do an experiment to measure the acceleration of Earth's gravity not on its surface, but twice as far from its center. The Earth's radius is about 6400 km, so let's do this 12,800 km from its center. What will you measure for the acceleration of gravity there?
Answer #1

acceleration a = Gm/r²

a is in m/s²

m is mass of earth

G = 6.67*10^-11 m^3kg/s^2

earth mass m= 5.974*10^24 kg

earth radius 6400 km

AT r = 12800 km
=> a = Gm/r²
a = (6.67*10^-11)*(5.974*10^24)/(12800*10^3)^2
a = 2.43 m/s^2